# Beyond Sparsity: Quantum Block Encoding for Dense Matrices via Hierarchically Low Rank Compression

### v1

@@ I Introduction
To solve linear systems in the quantum setting, we first review the definition of the quantum linear system problem (QLSP) [childsQuantumAlgorithmSystems2017]. In the following presentation, if not explicitly mentioned, we always denote $\|\cdot\|$ the 22-norm.
@@
Consider an $N\times N$ Hermitian matrix AA, which has condition number $\kappa$ and is normalized such that $\|A\|=1$. Let $\vec{b}$ be an NN-dimensional vector and $\vec{x}:=A^{-1}\vec{b}$. The corresponding quantum states are defined as
@@
An $(n+a)$-qubit unitary $U$ is an $(\alpha,a,\varepsilon)$-block-encoding of a $2^{n}\times 2^{n}$ matrix AA if it satisfies:
@@
The construction of block encodings can be achieved through several established methods, particularly for matrices with specific sparsity structures. We call a matrix AA is $s$-row-sparse ($s$-column-sparse) if each row (column) of AA has at most $s$ nonzero elements. The following lemma, adapted from Gilyén et al. [gilyenQuantumSingularValue2019] and summarized in Nguyen et al. [nguyenBlockencodingDenseFullrank2022], provides an example for oracle accessible sparse matrices.
@@
where $\tilde{a}_{ij}$ is the $b$-qubit fixed-point representation of $a_{ij}/\hat{a}$. Then, one can implement an $(\hat{a}\sqrt{s_{r}s_{c}},n+3,\varepsilon)$-block-encoding of AA with a single query to $\mathcal{O}_{r}$ and $\mathcal{O}_{c}$, two queries to $\mathcal{O}_{A}$, and $O(n+{\rm polylog}(\frac{\hat{a}\sqrt{s_{r}s_{c}}}{\varepsilon}))$ additional one and two qubit gates, using $O(b,{\rm polylog}(\frac{\hat{a}\sqrt{s_{r}s_{c}}}{\varepsilon}))$ ancilla qubits.
@@ III Hierarchically block separable matrices
The HBS matrix format provides an efficient way to represent this structure [hoFastDirectSolver2012, mindenRecursiveSkeletonizationFactorization2017]. To define this structure formally, consider a dense matrix $A\in\mathbb{C}^{N\times N}$. Let its row and column index sets be $I=\{1,\dots,N\}$ and $J=\{1,\dots,N\}$, respectively. We begin by partitioning the row index set II into $p$ disjoint blocks, $\{I_{j}\}_{j=1}^{p}$. The $j$-th row block, $I_{j}$, contains $n_{j}=|I_{j}|$ indices and is defined as:
@@
where $x_{j}\in\mathbb{C}^{n^{\prime}_{j}}$ and $b_{i}\in\mathbb{C}^{n_{i}}$ are the corresponding blocks of the vectors $x$ and $b$, and $A_{ij}\in\mathbb{C}^{n_{i}\times n^{\prime}_{j}}$ is the submatrix at the intersection of row block $I_{i}$ and column block $J_{j}$. Given an error tolerance $\varepsilon$, a matrix AA is defined as block separable if each of its off-diagonal blocks, $A_{ij}$ for $i\neq j$, admits a low-rank factorization of the form:
@@
Let $A\in\mathbb{C}^{m\times n}$ be a matrix with a target rank $k<\min(m,n)$. A rank-$k$ ID of AA is a factorization of the form $A\approx A_{\varepsilon}=BP$ under a given error tolerance $\varepsilon$, where $B\in\mathbb{C}^{m\times k}$ is a skeleton matrix whose columns are a specific subset of $k$ columns from AA. The matrix $P\in\mathbb{C}^{k\times n}$ is an interpolation matrix that contains the $k\times k$ identity matrix as a submatrix.
Efficient randomized algorithms can compute the ID with a typical complexity of $O(mn\log(k)+k^{2}n)$ [halkoFindingStructureRandomness2011]. ID is preferred over other low-rank methods like the singular value decomposition (SVD) for several reasons. Firstly, the skeleton matrix $B$ consists of actual columns from AA, which preserves physical structure. Secondly, the interpolation matrix $P$ has a provably bounded norm, ensuring numerical stability. We use the ID algorithm to approximate the off-diagonal blocks $A_{ij}$ (for $i\neq j$) as $A_{ij}\approx L_{i}S_{ij}R_{j}$, where $S_{ij}\in\mathbb{C}^{k_{i}^{r}\times k_{j}^{c}}$ with ranks satisfying $k_{i}^{r}\ll n_{i}$ and $k_{j}^{c}\ll n_{j}$.
@@
Figure 1: This figure illustrates the algorithmic process of the recursive HBS factorization. At level 1, (a) the matrix AA is decomposed into its near-field blocks ($D_{1}$) and far-field remainder ($B_{1}$). (b) The far-field part is approximated as $B_{1}\approx L_{1}S_{1}R_{1}$ using ID. (c) The process is recursively applied to $S_{1}$, and the indices of the resulting far-field part $B_{2}$ are reordered for the subsequent level.
@@
    The ranks of the off-diagonal blocks corresponding to well-separated points are bounded by a constant $r$ that depends on the desired precision but not on the matrix size NN. We therefore assume $k_{l}=r$ for all levels $l$.
@@
The classical procedure for constructing the decomposition of matrix AA leads to the following result regarding its computational cost and sparsity.
@@
The proof of Theorem III.1 can be found in Appendix LABEL:ap:classic. This theorem shows that the classical preprocessing cost and the number of non-zero elements in the resulting matrices both scale linearly with the matrix dimension NN. Furthermore, the entries of these matrices are bounded. These properties are important considerations for quantum algorithms, and we will apply the results of Theorem III.1 in the analysis that follows.
@@ IV A Sparsification Based Quantum Approach
The HBS decomposition yields factors $L,D,$ and $R$ that are either sparse or much smaller than the original matrix. This property holds at each level of the hierarchy, allowing us to approximate the original matrix AA with controlled accuracy. We begin by considering a single level of compression. In this case, the matrix AA is approximated as $A\approx D+LSR$. The linear system $Ax=b$ can then be solved by replacing AA with this approximation:
@@
We denote the block matrix in this system by $A_{\text{sp}}$ and refer to it as the extended sparsification of AA. This structure has been employed in [hoFastDirectSolver2012] for the construction of a fast direct solver for boundary integral equations.
To be used in a quantum algorithm, we must analyze the properties of $A_{\text{sp}}$. As shown in Theorem III.1, the total number of non-zero elements in $A_{\text{sp}}$ scales as $O(N)$. We now consider its row and column sparsity. The non-zero entries in any given row of $A_{\text{sp}}$ come from at most three types of block matrices: a diagonal block $D_{l}$, the interpolation matrices $L_{l}$ and $R_{l-1}$, and an identity matrix II. By construction, the interpolation matrices at level $l$ are sparse, with at most $k_{l}$ or $k_{l}+1$ non-zeros per row or column. The diagonal blocks $D_{l}$ contain near-neighbor interactions. At the finest level ($l=1$), a block interacts with its $3^{d}-1$ neighbors, so each row in $D_{1}$ has at most $3^{d}n_{1}$ non-zeros. For coarser levels ($l>1$), a block interacts with its $(6^{d}-3^{d})$ neighbors, resulting in at most $(6^{d}-3^{d})n_{l}$ non-zeros per row in $D_{l}$. Combining these bounds, the maximum number of non-zeros in any row of $A_{\text{sp}}$ is
@@
Finally, we consider the dimension of the matrix $A_{\text{sp}}$. The total number of rows, $N_{sp}$, is the sum of the dimensions of all the subblocks, which also scales linearly with NN:
@@
With these properties established, we can determine the cost of creating a block encoding for $A_{\text{sp}}$. Let $c_{\text{sp}}$ be the upper bound on the magnitude of elements in the original matrix AA, and let $f$ be the parameter defined in Theorem LABEL:apth:RRQR for the strong RRQR algorithm [guEfficientAlgorithmsComputing1996]. Based on the conditions of Theorem II.1, we arrive at the following theorem:
@@
Let AA be a matrix represented hierarchically to depth $\lambda$ with the extended sparsification $A_{\rm{sp}}=(a_{ij})_{N_{\rm{sp}}\times N_{\rm{sp}}}$ that is $s_{r}$-row-sparse and $s_{c}$-column-sparse. Then, one can implement an $(c_{\rm{sp}}\sqrt{s_{r}s_{c}},O(n),\varepsilon)$ block encoding of $A_{\rm{sp}}$ with a single query to $\mathcal{O}_{r}$ and $\mathcal{O}_{c}$, two queries to $\mathcal{O}_{A}$, and $O(n+{\rm polylog}(\frac{c_{\rm sp}\sqrt{s_{r}s_{c}}}{\varepsilon}))$ additional one and two qubit gates, using $O(b,{\rm polylog}(\frac{c_{\rm sp}\sqrt{s_{r}s_{c}}}{\varepsilon}))$ ancilla qubits. The oracles $O_{r},O_{c},O_{A}$ provide access to $O(N)$ nonzero matrix entries, where NN is the dimension of AA.
While this sparsification method allows us to handle the dense matrix with algorithms designed for sparse systems, it also introduces three challenges. First, the solution vector of the extended system contains not only the desired solution $x$ but also the auxiliary variables $y_{l}$ and $z_{l}$. These additional components reduce the probability of measuring $x$, which increases the overall runtime of the algorithm. Second, the complexity of quantum linear system solvers is sensitive to the condition number of the matrix. The relationship between the condition number of the extended matrix $A_{\text{sp}}$ and the original matrix AA is not obvious, and the transformation could potentially make it much larger. Finally, the HBS representation is an approximation of the original dense matrix, and the impact of this initial error on the accuracy of the final quantum solution needs to be analyzed. Theorem IV.2 provides quantitative bounds for each of these issues.
@@ V A Direct Block Encoding Algorithm for Hierarchical Matrices
The construction relies on two main quantum primitives: the linear combination of unitaries (LCU) and quantum matrix multiplication (QMM). See Figures LABEL:fig:lcu_circuit and LABEL:fig:qmm_circuit for their quantum circuits, respectively. We assume that block encodings for the component matrices $D_{l}$, $L_{l}$, and $R_{l}$ are given, denoted as $U^{D}_{l}$, $U^{L}_{l}$, and $U_{l}^{R}$, respectively. Each encoding uses $a$ ancilla qubits and has an error of $\varepsilon$. The subnormalization factors for the diagonal blocks $D_{l}$ are denoted $\alpha_{l}^{(D)}$, and for the interpolation matrices $L_{l},R_{l}$ they are $\alpha_{l}^{(L)}$. Based on Lemma II.1 for block encoding sparse matrices, and recalling that the elements of $D_{l}$ are drawn from AA, we have

### v2

@@ I Introduction
To solve linear systems in the quantum setting, we first review the definition of the quantum linear system problem (QLSP) [childsQuantumAlgorithmSystems2017]. In the following presentation, if not explicitly mentioned, we always denote $\|\cdot\|$ the $2$-norm.
@@
Consider an $N\times N$ Hermitian matrix $A$, which has condition number $\kappa$ and is normalized such that $\|A\|=1$. Let $\vec{b}$ be an $N$-dimensional vector and $\vec{x}:=A^{-1}\vec{b}$. The corresponding quantum states are defined as
@@
An $(n+a)$-qubit unitary $U$ is an $(\alpha,a,\varepsilon)$-block-encoding of a $2^{n}\times 2^{n}$ matrix $A$ if it satisfies:
@@
The construction of block encodings can be achieved through several established methods, particularly for matrices with specific sparsity structures. We call a matrix $A$ is $s$-row-sparse ($s$-column-sparse) if each row (column) of $A$ has at most $s$ nonzero elements. The following lemma, adapted from Gilyén et al. [gilyenQuantumSingularValue2019] and summarized in Nguyen et al. [nguyenBlockencodingDenseFullrank2022], provides an example for oracle accessible sparse matrices.
@@
where $\tilde{a}_{ij}$ is the $b$-qubit fixed-point representation of $a_{ij}/\hat{a}$. Then, one can implement an $(\hat{a}\sqrt{s_{r}s_{c}},n+3,\varepsilon)$-block-encoding of $A$ with a single query to $\mathcal{O}_{r}$ and $\mathcal{O}_{c}$, two queries to $\mathcal{O}_{A}$, and $O(n+{\rm polylog}(\frac{\hat{a}\sqrt{s_{r}s_{c}}}{\varepsilon}))$ additional one and two qubit gates, using $O(b,{\rm polylog}(\frac{\hat{a}\sqrt{s_{r}s_{c}}}{\varepsilon}))$ ancilla qubits.
@@ III Hierarchically block separable matrices
The HBS matrix format provides an efficient way to represent this structure [hoFastDirectSolver2012, mindenRecursiveSkeletonizationFactorization2017]. To define this structure formally, consider a dense matrix $A\in\mathbb{C}^{N\times N}$. Let its row and column index sets be $I=\{1,\dots,N\}$ and $J=\{1,\dots,N\}$, respectively. We begin by partitioning the row index set $I$ into $p$ disjoint blocks, $\{I_{j}\}_{j=1}^{p}$. The $j$-th row block, $I_{j}$, contains $n_{j}=|I_{j}|$ indices and is defined as:
@@
where $x_{j}\in\mathbb{C}^{n^{\prime}_{j}}$ and $b_{i}\in\mathbb{C}^{n_{i}}$ are the corresponding blocks of the vectors $x$ and $b$, and $A_{ij}\in\mathbb{C}^{n_{i}\times n^{\prime}_{j}}$ is the submatrix at the intersection of row block $I_{i}$ and column block $J_{j}$. Given an error tolerance $\varepsilon$, a matrix $A$ is defined as block separable if each of its off-diagonal blocks, $A_{ij}$ for $i\neq j$, admits a low-rank factorization of the form:
@@
Let $A\in\mathbb{C}^{m\times n}$ be a matrix with a target rank $k<\min(m,n)$. A rank-$k$ ID of $A$ is a factorization of the form $A\approx A_{\varepsilon}=BP$ under a given error tolerance $\varepsilon$, where $B\in\mathbb{C}^{m\times k}$ is a skeleton matrix whose columns are a specific subset of $k$ columns from $A$. The matrix $P\in\mathbb{C}^{k\times n}$ is an interpolation matrix that contains the $k\times k$ identity matrix as a submatrix.
Efficient randomized algorithms can compute the ID with a typical complexity of $O(mn\log(k)+k^{2}n)$ [halkoFindingStructureRandomness2011]. ID is preferred over other low-rank methods like the singular value decomposition (SVD) for several reasons. Firstly, the skeleton matrix $B$ consists of actual columns from $A$, which preserves physical structure. Secondly, the interpolation matrix $P$ has a provably bounded norm, ensuring numerical stability. We use the ID algorithm to approximate the off-diagonal blocks $A_{ij}$ (for $i\neq j$) as $A_{ij}\approx L_{i}S_{ij}R_{j}$, where $S_{ij}\in\mathbb{C}^{k_{i}^{r}\times k_{j}^{c}}$ with ranks satisfying $k_{i}^{r}\ll n_{i}$ and $k_{j}^{c}\ll n_{j}$.
@@
Figure 1: This figure illustrates the algorithmic process of the recursive HBS factorization. At level 1, (a) the matrix $A$ is decomposed into its near-field blocks ($D_{1}$) and far-field remainder ($B_{1}$). (b) The far-field part is approximated as $B_{1}\approx L_{1}S_{1}R_{1}$ using ID. (c) The process is recursively applied to $S_{1}$, and the indices of the resulting far-field part $B_{2}$ are reordered for the subsequent level.
@@
    The ranks of the off-diagonal blocks corresponding to well-separated points are bounded by a constant $r$ that depends on the desired precision but not on the matrix size $N$. We therefore assume $k_{l}=r$ for all levels $l$.
@@
The classical procedure for constructing the decomposition of matrix $A$ leads to the following result regarding its computational cost and sparsity.
@@
The proof of Theorem III.1 can be found in Appendix LABEL:ap:classic. This theorem shows that the classical preprocessing cost and the number of non-zero elements in the resulting matrices both scale linearly with the matrix dimension $N$. Furthermore, the entries of these matrices are bounded. These properties are important considerations for quantum algorithms, and we will apply the results of Theorem III.1 in the analysis that follows.
@@ IV A Sparsification Based Quantum Approach
The HBS decomposition yields factors $L,D,$ and $R$ that are either sparse or much smaller than the original matrix. This property holds at each level of the hierarchy, allowing us to approximate the original matrix $A$ with controlled accuracy. We begin by considering a single level of compression. In this case, the matrix $A$ is approximated as $A\approx D+LSR$. The linear system $Ax=b$ can then be solved by replacing $A$ with this approximation:
@@
We denote the block matrix in this system by $A_{\text{sp}}$ and refer to it as the extended sparsification of $A$. This structure has been employed in [hoFastDirectSolver2012] for the construction of a fast direct solver for boundary integral equations.
To be used in a quantum algorithm, we must analyze the properties of $A_{\text{sp}}$. As shown in Theorem III.1, the total number of non-zero elements in $A_{\text{sp}}$ scales as $O(N)$. We now consider its row and column sparsity. The non-zero entries in any given row of $A_{\text{sp}}$ come from at most three types of block matrices: a diagonal block $D_{l}$, the interpolation matrices $L_{l}$ and $R_{l-1}$, and an identity matrix $I$. By construction, the interpolation matrices at level $l$ are sparse, with at most $k_{l}$ or $k_{l}+1$ non-zeros per row or column. The diagonal blocks $D_{l}$ contain near-neighbor interactions. At the finest level ($l=1$), a block interacts with its $3^{d}-1$ neighbors, so each row in $D_{1}$ has at most $3^{d}n_{1}$ non-zeros. For coarser levels ($l>1$), a block interacts with its $(6^{d}-3^{d})$ neighbors, resulting in at most $(6^{d}-3^{d})n_{l}$ non-zeros per row in $D_{l}$. Combining these bounds, the maximum number of non-zeros in any row of $A_{\text{sp}}$ is
@@
Finally, we consider the dimension of the matrix $A_{\text{sp}}$. The total number of rows, $N_{sp}$, is the sum of the dimensions of all the subblocks, which also scales linearly with $N$:
@@
With these properties established, we can determine the cost of creating a block encoding for $A_{\text{sp}}$. Let $c_{\text{sp}}$ be the upper bound on the magnitude of elements in the original matrix $A$, and let $f$ be the parameter defined in Theorem LABEL:apth:RRQR for the strong RRQR algorithm [guEfficientAlgorithmsComputing1996]. Based on the conditions of Theorem II.1, we arrive at the following theorem:
@@
Let $A$ be a matrix represented hierarchically to depth $\lambda$ with the extended sparsification $A_{\rm{sp}}=(a_{ij})_{N_{\rm{sp}}\times N_{\rm{sp}}}$ that is $s_{r}$-row-sparse and $s_{c}$-column-sparse. Then, one can implement an $(c_{\rm{sp}}\sqrt{s_{r}s_{c}},O(n),\varepsilon)$ block encoding of $A_{\rm{sp}}$ with a single query to $\mathcal{O}_{r}$ and $\mathcal{O}_{c}$, two queries to $\mathcal{O}_{A}$, and $O(n+{\rm polylog}(\frac{c_{\rm sp}\sqrt{s_{r}s_{c}}}{\varepsilon}))$ additional one and two qubit gates, using $O(b,{\rm polylog}(\frac{c_{\rm sp}\sqrt{s_{r}s_{c}}}{\varepsilon}))$ ancilla qubits. The oracles $O_{r},O_{c},O_{A}$ provide access to $O(N)$ nonzero matrix entries, where $N$ is the dimension of $A$.
While this sparsification method allows us to handle the dense matrix with algorithms designed for sparse systems, it also introduces three challenges. First, the solution vector of the extended system contains not only the desired solution $x$ but also the auxiliary variables $y_{l}$ and $z_{l}$. These additional components reduce the probability of measuring $x$, which increases the overall runtime of the algorithm. Second, the complexity of quantum linear system solvers is sensitive to the condition number of the matrix. The relationship between the condition number of the extended matrix $A_{\text{sp}}$ and the original matrix $A$ is not obvious, and the transformation could potentially make it much larger. Finally, the HBS representation is an approximation of the original dense matrix, and the impact of this initial error on the accuracy of the final quantum solution needs to be analyzed. Theorem IV.2 provides quantitative bounds for each of these issues.
@@ V A Direct Block Encoding Algorithm for Hierarchical Matrices
The construction relies on two main quantum primitives: the linear combination of unitaries (LCU) and quantum matrix multiplication (QMM). See Figures LABEL:fig:lcu_circuit and LABEL:fig:qmm_circuit for their quantum circuits, respectively. We assume that block encodings for the component matrices $D_{l}$, $L_{l}$, and $R_{l}$ are given, denoted as $U^{D}_{l}$, $U^{L}_{l}$, and $U_{l}^{R}$, respectively. Each encoding uses $a$ ancilla qubits and has an error of $\varepsilon$. The subnormalization factors for the diagonal blocks $D_{l}$ are denoted $\alpha_{l}^{(D)}$, and for the interpolation matrices $L_{l},R_{l}$ they are $\alpha_{l}^{(L)}$. Based on Lemma II.1 for block encoding sparse matrices, and recalling that the elements of $D_{l}$ are drawn from $A$, we have